OpenStudy (anonymous):

Deriving the equation of a parabola, (x-4)^2+y^2-16y+16=y^2+4y^2+4 Been stuck on this forever.

OpenStudy (anonymous):

what on earth is that?

OpenStudy (anonymous):

I'm supposed to simplify it

OpenStudy (anonymous):

\[(x-4)^2+y^2-16y+16=y^2+4y^2+4 \] combine like terms on the right and get \[(x-4)^2+y^2-16y+16=5y^2+4 \]

OpenStudy (anonymous):

then subtract if you want to set it equal to zero and get \[(x-4)^2-4y^2-16y+12=0\] if you like it is not a parabola for sure

OpenStudy (anonymous):

really

OpenStudy (anonymous):

it is a hyperbola was that the original question?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

ok so

OpenStudy (anonymous):

it started off as two distance formulas

OpenStudy (anonymous):

im try to find the distance between the directx and the focus

OpenStudy (anonymous):

the directx being 0,2

OpenStudy (anonymous):

the focus being 4,4

OpenStudy (anonymous):

the directrix of a parabola is not a point, it is a line

OpenStudy (anonymous):

I know the answer but i need to show my work and the problem says to set up the problem as two distance formulas being equal to each other and to simplify the equation

OpenStudy (anonymous):

is it the line \(y=2\)? it cannot be a point \((0,2)\) although that could be a point on the line

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the distance between the line \(y=2\) and the point \((4,4)\) does not use the distance formula that distance is \(2\)

OpenStudy (anonymous):

i mean -2

OpenStudy (anonymous):

not 2

OpenStudy (anonymous):



OpenStudy (anonymous):

how far down from the point \((4,4)\) is the line \(y=-2\)?

OpenStudy (anonymous):

6 units right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lesson 9.05 page 3

hero (hero):

The link doesn't work, but you can represent the directrix as a point if you write it as (x , 2)

OpenStudy (anonymous):

Ok so im supposed to explain how to find the equation of the parabola given the focal point and the directrix.

OpenStudy (anonymous):

\[4p(y-k)^2=(x-h)^2\] where the vertex is \((h,k)\) and the distance between the focus and then vertex (which is equal to the distance between the vertex and the directrix, which is half the distance between the focus and the directix) is \(p\)

OpenStudy (anonymous):

that was wrong sorry it is \[4p(y-k)=(x-h)^2\]

OpenStudy (anonymous):

so if the focus is \((4,4)\) and the directrix is \(y=-2\) then the vertex is \((4,1)\) half way between them

OpenStudy (anonymous):

the distance between \((4,1)\) and \((4,4)\) is \(3\) so the equation is \[4\times 3(y-1)=(x-4)^2\]

OpenStudy (anonymous):

here is the check note the focus and the directrix http://www.wolframalpha.com/input/?i=parabola+12%28y-1%29%3D%28x-4%29^2

hero (hero):

@satellite73, @EternalRaze said earlier that the problem says to set up the problem as two distance formulas being equal to each other and to simplify the equation.

OpenStudy (anonymous):

ick no one used the distance formula to find a vertical distance that is using pythagoras, but vertical distance doesn't use it the distance between \((a, b)\) and \((a,c)\) is \(|b-c|\)

hero (hero):

@satellite73, if someone told me I had to solve x^2 + 5x + 6 using the quadratic formula, I would have to use the quadratic formula rather than factoring even though I prefer factoring better.